# Data-Driven Approach to Capitation Reform in Rwanda

### v1

@@ Abstract
As part of Rwanda’s transition toward universal health coverage, the national Community-Based Health Insurance (CBHI) scheme is moving from retrospective fee-for-service reimbursements to prospective capitation payments for public primary healthcare providers. This report outlines a data-driven approach to designing, calibrating, and monitoring the capitation model using individual-level claims data from the Intelligent Health Benefits System (IHBS). We introduce a transparent, interpretable formula for allocating payments to Health Centers and their affiliated Health Posts. The formula is based on catchment population, service utilization patterns, and patient inflows, with parameters estimated via regression models calibrated on national claims data. Repeated validation exercises show the payment scheme closely aligns with historical spending while promoting fairness and adaptability across diverse facilities. In addition to payment design, the same dataset enables actionable behavioral insights. We highlight the use case of monitoring antibiotic prescribing patterns, particularly in pediatric care, to flag potential overuse and guideline deviations. Together, these capabilities lay the groundwork for a learning health financing system: one that connects digital infrastructure, resource allocation, and service quality to support continuous improvement and evidence-informed policy reform.
@@ 1 Introduction
This report documents the implementation of Rwanda’s national capitation model for CBHI-financed primary care. It presents:
@@ 2 Primary Healthcare Delivery and Financing in Rwanda
Under capitation, each Health Center will be responsible for delivering a standardized package of primary health care services to the population within its designated catchment area. These catchments are defined at the village level and assigned by the Ministry of Health (MoH), although CBHI beneficiaries may also seek care outside their assigned catchment. Some Health Centers are designated as Medicalized Health Centers because they provide additional services, such as cesarean sections, that fall outside the standard PHC package. These services are reimbursed separately through FFS and are excluded from the capitation model described in this report.
@@ 3 The Capitation Reform
Data-driven capitation design is enabled by the rich data captured through IHBS, which allows for capitation payments grounded in real-world utilization patterns and facility-level monitoring, dynamic adjustment, and policy feedback loops. All analyses documented in this report were performed on anonymized datasets in compliance with national data protection regulations, preserving confidentiality while ensuring modeling fidelity.
@@
These principles shaped a methodology rooted in empirical data, transparency, and operational realism, ensuring that the resulting capitation model aligns financial incentives with health system goals while maintaining provider engagement and service quality. Methodological development combined data engineering, descriptive analytics, and parameter estimation. The dataset used for the analyses presented in this report spanned April 1, 2023 to March 4, 2025, and included visit-level and membership data from IHBS, enriched with household location and catchment definitions.
@@ 3.4 Estimated Capitation Payments
The estimated parameters using data for January to December 2024 (as compared to the fiscal year used in this report’s analysis) yield 867, 1225, and 1506 for the low, medium, and high utilization $A$ parameters and 1142 for the $B$ inflow component, i.e., all variations with respect to the 12-month period six months earlier are below 5%, indicating the estimation is relatively stable while still showing a trend of diminishing average cost for within catchment visits while no variation to slight increment in the inflow visits.
@@ 3.6 Robustness of the Capitation Formula and Alternative Approaches
Several alternative approaches were explored during the design of the capitation formula, though the results are not described in this report. One involved modifying the inflow component using a binary indicator for Health Centers located within 2 km of a hospital. These facilities often act as referral gatekeepers, attracting low-cost visits from outside their catchment area. While performance improved slightly, further analysis revealed no consistent patterns in cost per visit or referral rates among these centers, so this approach was discarded due to potential implementation complexity.
@@ 4 Data-Driven Insights for Behavioral Changes: Antibiotic Use
To explore potential quality issues in prescribing practices in Rwanda, we conducted an analysis of consultations for children under 15 years of age. We focused on a subset of pediatric visits in which all diagnoses within a given consultation belonged exclusively to a single diagnosis category, such as lower respiratory infections or gastrointestinal conditions (different consultations could fall into different categories). This restriction allowed for more clinically coherent comparisons across visits. The results presented in this report are restricted to visits to Health Posts in 2024; however, similar results are obtained when analyzing visits to Health Centers.
@@ 5 Conclusions
This report describes a methodology for the implementation and iterative refinemnt of a national-scale capitation system using high-quality health data. Rwanda’s IHBS, now deployed across all public primary care facilities, enables a dynamic, data-driven financing model grounded in real-world utilization patterns. The capitation formula promotes equitable budget allocation across Health Centers and their affiliated Health Posts by accounting for intra- and inter-catchment patient flows, thereby reflecting patient mobility. Its design combines a regression-based approach with quarterly payment estimation using prior-year data to account for seasonal variations in demand. A prospective adjustment mechanism further refines fairness by correcting payments for facilities experiencing significant shifts in utilization. Early results suggest the model aligns well with historical costs for most facilities, simplifies administrative processes, and lays a foundation for quality-focused resource allocation. It is important to note that while historical reimbursement patterns are used here as a benchmark to assess the accuracy of the capitation model, they are not necessarily an ideal standard. The goal of capitation is to shift provider incentives toward efficiency and quality; some divergence from historical costs may reflect progress toward these objectives rather than a model limitation. Transparent and scalable, this framework advances Rwanda’s universal health coverage goals and could offer a replicable model for other low- and middle-income countries transitioning to data-informed health financing.

### v2

@@ Abstract
As part of Rwanda’s transition toward universal health coverage, the national Community-Based Health Insurance (CBHI) scheme is moving from retrospective fee-for-service reimbursements to prospective capitation payments for public primary healthcare providers. This work outlines a data-driven approach to designing, calibrating, and monitoring the capitation model using individual-level claims data from the Intelligent Health Benefits System (IHBS). We introduce a transparent, interpretable formula for allocating payments to Health Centers and their affiliated Health Posts. The formula is based on catchment population, service utilization patterns, and patient inflows, with parameters estimated via regression models calibrated on national claims data. Repeated validation exercises show the payment scheme closely aligns with historical spending while promoting fairness and adaptability across diverse facilities. In addition to payment design, the same dataset enables actionable behavioral insights. We highlight the use case of monitoring antibiotic prescribing patterns, particularly in pediatric care, to flag potential overuse and guideline deviations. Together, these capabilities lay the groundwork for a learning health financing system: one that connects digital infrastructure, resource allocation, and service quality to support continuous improvement and evidence-informed policy reform.
@@ 1 Introduction
This work documents the implementation of Rwanda’s national capitation model for CBHI-financed primary care. It presents:
@@ 2 Primary Healthcare Delivery and Financing in Rwanda
Under capitation, each Health Center will be responsible for delivering a standardized package of primary health care services to the population within its designated catchment area. These catchments are defined at the village level and assigned by the Ministry of Health (MoH), although CBHI beneficiaries may also seek care outside their assigned catchment. Some Health Centers are designated as Medicalized Health Centers because they provide additional services, such as cesarean sections, that fall outside the standard PHC package. These services are reimbursed separately through FFS and are excluded from the capitation model described in this article.
@@ 3 The Capitation Reform
Data-driven capitation design is enabled by the rich data captured through IHBS, which allows for capitation payments grounded in real-world utilization patterns and facility-level monitoring, dynamic adjustment, and policy feedback loops. All analyses documented in this work were performed on anonymized datasets in compliance with national data protection regulations, preserving confidentiality while ensuring modeling fidelity.
@@
These principles shaped a methodology rooted in empirical data, transparency, and operational realism, ensuring that the resulting capitation model aligns financial incentives with health system goals while maintaining provider engagement and service quality. Methodological development combined data engineering, descriptive analytics, and parameter estimation. The dataset used for the analyses presented in this work spanned April 1, 2023 to March 4, 2025, and included visit-level and membership data from IHBS, enriched with household location and catchment definitions.
@@ 3.4 Estimated Capitation Payments
The estimated parameters using data for January to December 2024 (as compared to the fiscal year used in this work’s analysis) yield 867, 1225, and 1506 for the low, medium, and high utilization $A$ parameters and 1142 for the $B$ inflow component, i.e., all variations with respect to the 12-month period six months earlier are below 5%, indicating the estimation is relatively stable while still showing a trend of diminishing average cost for within catchment visits while no variation to slight increment in the inflow visits.
@@ 3.6 Robustness of the Capitation Formula and Alternative Approaches
Several alternative approaches were explored during the design of the capitation formula, though the results are not described in this work. One involved modifying the inflow component using a binary indicator for Health Centers located within 2 km of a hospital. These facilities often act as referral gatekeepers, attracting low-cost visits from outside their catchment area. While performance improved slightly, further analysis revealed no consistent patterns in cost per visit or referral rates among these centers, so this approach was discarded due to potential implementation complexity.
@@ 4 Data-Driven Insights for Behavioral Changes: Antibiotic Use
To explore potential quality issues in prescribing practices in Rwanda, we conducted an analysis of consultations for children under 15 years of age. We focused on a subset of pediatric visits in which all diagnoses within a given consultation belonged exclusively to a single diagnosis category, such as lower respiratory infections or gastrointestinal conditions (different consultations could fall into different categories). This restriction allowed for more clinically coherent comparisons across visits. The results presented in this article are restricted to visits to Health Posts in 2024; however, similar results are obtained when analyzing visits to Health Centers.
@@ 5 Conclusions
This article describes a methodology for the implementation and iterative refinemnt of a national-scale capitation system using high-quality health data. Rwanda’s IHBS, now deployed across all public primary care facilities, enables a dynamic, data-driven financing model grounded in real-world utilization patterns. The capitation formula promotes equitable budget allocation across Health Centers and their affiliated Health Posts by accounting for intra- and inter-catchment patient flows, thereby reflecting patient mobility. Its design combines a regression-based approach with quarterly payment estimation using prior-year data to account for seasonal variations in demand. A prospective adjustment mechanism further refines fairness by correcting payments for facilities experiencing significant shifts in utilization. Early results suggest the model aligns well with historical costs for most facilities, simplifies administrative processes, and lays a foundation for quality-focused resource allocation. It is important to note that while historical reimbursement patterns are used here as a benchmark to assess the accuracy of the capitation model, they are not necessarily an ideal standard. The goal of capitation is to shift provider incentives toward efficiency and quality; some divergence from historical costs may reflect progress toward these objectives rather than a model limitation. Transparent and scalable, this framework advances Rwanda’s universal health coverage goals and could offer a replicable model for other low- and middle-income countries transitioning to data-informed health financing.